OpenStudy (anonymous):

Calvin bought 5 packets of cheese and 3 burgers for $29.99. At the same prices, Alex bought 3 packets of cheese and 7 burgers for $32.71. Find the price of one packet of cheese by writing a system of equations, then solve using Cramer's Rule.

OpenStudy (anonymous):

Let \(x\) be the price of a packet of cheese and \(y\) the price of a burger. So we have the following system of equations: $$ \begin{cases} 5x + 3y = 29.99 \\ 3x + 7y = 32.71 \end{cases} $$ Can you find \(x\) and \(y\)?

OpenStudy (anonymous):

ummm im not sure how to find x and y :( can you help...

OpenStudy (anonymous):

ok, you can for example take the first equation $$ 5x + 3y = 29.99 $$ Now let's find 'x' in that: $$ 5x = 29.99 - 3y\\ x = \frac{29.99 - 3y}{5} $$ Now you can take value for 'x' and substitute with the 'x' in the second equation. Then you'll have an equation with single variable 'y' =] Once you find the numerical value of 'y' you can substitute it in the first equation and find such numerical value for 'x'. Clear enough?

OpenStudy (anonymous):

I don't know how to solve the equation you gave me though..

OpenStudy (anonymous):

x=29.99−3y 5 this equation..

OpenStudy (anonymous):

that's ok. the value of 'x', by the first equation, has to be equal to: $$ x = \frac{29.99 -3y}{5} $$ We can't get it any better than that. But, we can take this value and put it in the second equation: $$ 3x + 7y = 32.71 \\ 3\cdot\bigg( \frac{29.99 -3y}{5} \bigg) + 7y = 32.71 $$ Can you solve this one for 'y'?

OpenStudy (anonymous):

so.. if I solve the distributive form.. i get 89.97 - 9y over 15 right?

OpenStudy (anonymous):

and.. then..i divide 89.97 with 15? and get 5.998..

OpenStudy (anonymous):

eh, why divide by 15? we get (89.97 - 9y) / 5 for first term

OpenStudy (anonymous):

oh..bhhaha ok..:/

OpenStudy (anonymous):

and..do I have to add the two ys? omgsh. i have no idea.

OpenStudy (anonymous):

let's now multiply this by 5 to get rid of the factorial we should now now $$ \cancel 5 \cdot \bigg(\frac{89.97 - 9y}{\cancel5}\bigg) + 5 \cdot 7y = 5 \cdot 32.71 $$

OpenStudy (anonymous):

Can you find 'y' now?

OpenStudy (anonymous):

yea i think so :D wait,,,,

OpenStudy (anonymous):

um first i multiplied 5 times 32. 71 and I got 163.55

OpenStudy (anonymous):

then I have to add the two numbers without a y next to it right..?

OpenStudy (anonymous):

ok so far =)

OpenStudy (anonymous):

well ye, move and switch the sign. (basically subtracting 89.97 from both sides..)

OpenStudy (anonymous):

oh so i subtract 89.97 with 163.55 which is 735.8

OpenStudy (anonymous):

73.58 ye

OpenStudy (anonymous):

oh oops yeah. and do i move the five over too?

OpenStudy (anonymous):

no. we have 5*7y. which means 35*y, right?

OpenStudy (anonymous):

we need to add to that the remaining (-9*y) from the factorial we had.

OpenStudy (anonymous):

to clarify what we have now is basically: $$ 89.97 - 9y + 5\cdot7y = 163.55 \\ - 9y + 5\cdot7y = 163.55 - 89.97 \\ -9y + 35y = 73.58 $$

OpenStudy (anonymous):

-9y + 35y becomes what?

OpenStudy (anonymous):

26y

OpenStudy (anonymous):

so the answer is 2.83

OpenStudy (anonymous):

yes. now let's take this and put the new value of 'y' into the first equation $$ 5x + 3y = 29.99 $$

OpenStudy (anonymous):

so.. x is 4.3?

OpenStudy (anonymous):

yes here, same solution as wolfram alpha =] http://www.wolframalpha.com/input/?i=5x%2B3y+%3D+29.99%2C+3x%2B7y+%3D+32.71

OpenStudy (anonymous):

thank you so much :D

OpenStudy (anonymous):

well, half way. after I read that your question asks to solve it... I've seen that they also want you to use Cramer's Rule doing so..

OpenStudy (anonymous):

oh... ok

OpenStudy (anonymous):

$5.30 $1.83 $4.30 $2.83 there are four answers though

OpenStudy (anonymous):

so wouldn't it be the third one?

OpenStudy (anonymous):

Solving as we did is the 'simple' way for solving system of equations. you should be familiar with it first. If you only need the solution then you have them already. x was packet of cheese and y was burger. If you also need the way... then we can use Cramer's Rule.

OpenStudy (anonymous):

I think I like the simple way better. thanks though :D

OpenStudy (anonymous):

To be honest, I like it better too heh =) well you're welcome, good luck with the rest

OpenStudy (anonymous):

thanksss :DD